A health clinic has a solution of bleach to sterilize petri dishes in which culture are grown. the sterilization tank contains 100 gal of a solution of 2% ordinary household bleach mixed with pure distilled water. now research indicates that concentration of bleach should be 5% for complete sterilization. how much of the solution should be drain and replaced with bleach intent to the recommended level?
Answer by stanbon(75887)

A health clinic has a solution of bleach to sterilize petri dishes in which culture are grown. the sterilization tank contains 100 gal of a solution of 2% ordinary household bleach mixed with pure distilled water. now research indicates that concentration of bleach should be 5% for complete sterilization. how much of the solution should be drain and replaced with bleach intent to the recommended level?
------
Equation:
active - active + active = active
----
0.02*100 - 0.02x + 1.00*x = 0.05*100
----
Multiply thru by 100 to get:
2*100 - 2x + 100x = 5*100
---
98x = 3*100
x = 3.0612 gallons of 2% liquid should be replaced by 100% bleach.
=======================
Cheers,
Stan H.
=======================